# Towards Identifiability of Interventional Stochastic Differential Equations

### v3

@@ 1 Introduction
However, in many natural settings, data is modeled as following and SDE even if one does not have access to explicit trajectories. Studies of ecological systems focus on the long-term survival of multiple species modeled by the quasi-stationary state of SDEs with environmental factors as perturbations (Hening & Li, 2021). The application of flow cytometry to protein signaling networks under perturbation (Sachs et al., 2005) is destructive and yields protein quantification at one time point, modeled using the stationary distributions of linear SDEs in Varando & Hansen (2020).
One highly motivating application is single-cell genomic sequencing with high-throughput CRISPR perturbations. Biologists are often interested in inferring the gene regulatory network (GRN) that characterizes the dynamics of gene expression, informing which genes should be targeted for treatment (Dixit et al., 2016). But the destructive nature of sequencing makes it impossible to observe the trajectory of a single cell at multiple time-points, and in general it is difficult to obtain any time-series genomics data due to the high expense. Therefore, practitioners often only collect data at the end of an experiment, i.e., from the stationary distribution of the system.
Understanding the dynamics is essential for extrapolating to unseen settings, but noise and latent confounding makes it non-trivial to determine the true dynamics. Causal disentanglement aims to learn causal factors in spite of these confounders, mainly focusing on directed acyclic graph (DAG) based methods. To demonstrate these methods are well-founded, there is considerable effort devoted to understanding which models have identifiability guarantees (Lachapelle et al., 2022). However, these models suffer from inherent weakness, in particular 1) being unable to represent cycles or 2) approximate continuous-time dynamical models.
There has been renewed interest in modeling with stochastic differential equations (SDE) directly (Peters et al., 2022). In the genomic context, there is precedent for this type of modeling to represent the so- called “Waddington landscape” (Waddington, 2014), the hypothetical energy surface of cells. Furthermore, SDEs are commonly used for simulating transcriptomic datasets from a given gene regulatory network (Pratapa et al., 2020; Dibaeinia & Sinha, 2020).
As demonstrated in the context of diffusion models, SDEs are fully expressive in practice and can accurately generate observational data (Song et al., 2021). But to identify the true underlying SDE requires assumptions on the model. Foundational theoretical works on identifying dynamical systems typically learn from many trajectories or even the infinitesimal generator of the dynamics (Hansen & Sokol, 2014), leaving open the harder setting of observing only the stationary distribution.
@@ 2.2 Stochastic Differential Equations
We will only consider autonomous systems, i.e. where the drift and noise terms have no dependence on time $t$. We will enforce the weak conditions on the drift and noise to guarantee a unique stationary distribution (Berglund, 2021), with a density $p$ that satisfies the Fokker-Planck equation,
@@ 2.4 Interventional SDEs
We denote the concatenated intervention column vectors by the matrix $C\in\mathbb{R}^{n\times k}$. In the causal disentanglement literature, shift interventions typically give fewer guarantees (Buchholz et al., 2024; Squires et al., 2023). Therefore, we assume knowledge of the interventions:
@@ 3.1 Causal Representation
In terms of modeling causality (Pearl, 2009), the most popular underlying model is the structural causal model (SCM), which characterizes the conditional distribution of a random variable under arbitrary intervention. Learning an SCM typically requires very strong assumptions such as sparsity (Schölkopf et al., 2021), interventional data (Lachapelle et al., 2022), parametric assumptions (Peters & Bühlmann, 2014), among many other results. Sparsity is a very common theme in these models, though it may also be expressed in an assumption that the number of latent variables is small, i.e. a low-rank constraint (Fang et al., 2023).
@@ 3.2 Causal Disentanglement with Interventional Data
The bulk of the literature on causal disentanglement focuses on SCMs with an underlying DAG. Relevant to our work are results that assume access to multiple interventional environments, either acting directly on the observed variables (Brouillard et al., 2020) or identifying a latent model under some distributional assumptions (Lachapelle et al., 2022; Squires et al., 2023; Buchholz et al., 2024).
Some of these works have addressed the crucial limitation of acyclicity (Zheng et al., 2018; Lee et al., 2019; Atanackovic et al., 2023), but without necessarily incorporating dynamics. Many works require hard interventions, where an intervened variable is a function of exogenous noise, although some can handle soft interventions (Zhang et al., 2024).
@@ 3.3 Dynamical System Methods
Previous work has considered modeling perturbations with dynamical systems (Peters et al., 2022). One can prove identifiability of SDE parameters from the generator (Hansen & Sokol, 2014) or trajectories (Guan et al., 2024). Other methods work in our harder setting where observed data is drawn from the stationary distribution under an SDE, and match a learned SDE under numerous interventions. These works focus on linear drift and intervention-dependent parameters (Rohbeck et al., 2024) or non-linear drift with shift interventions (Lorch et al., 2024). Notably, neither paper gives theory to confirm if these models are identifiable.
There are also methods specific to a particular scientific domain. In genomics, some methods act on pseudotime, an inferred notion of time from cell states when very few “real” timepoints are available (Wang et al., 2024; Hossain et al., 2024). Although harder to obtain due the destructive nature of RNA sequencing, genuine temporal data (even with very few timepoints) can also be modeled with the intent of extracting a GRN (Lin et al., 2025).
The closest work to ours proves an identifiability bound for linear SDEs under a strong sparsity assumption (Dettling et al., 2023). However, their result doesn’t consider any interventional data, and the exact pattern of sparsity in the drift matrix is assumed to be known a priori, which is rarely the case in applied problems of interest. Another closely related work is Guan et al. (2024), which can simultaneously infer the drift and diffusion. However, this work applies only in an easier setting that assumes linearity, doesn’t study interventions, and assumes multiple temporal marginals rather than just the stationary distribution.
@@ 4 Main Results
We consider two main parameterizations of the drift as linear or a two-layer neural network (an MLP) to verify when the model may be uniquely identified up to appropriate invariances. Prior work has considered sparsity in the linear setting only (Dettling et al., 2023), but we focus on parameterizations that project to a low-dimensional space. In other words, although the ambient dimension $n$ may be large, we assume the hidden dimension $r$ is much smaller, at least $n>2r$, and ideally the number of interventions to uniquely identify the parameters should scale with $r$.
@@ 4.1 Linear case
We start with the linear case and consider the parameterization v(x)=(ABD)xv(x)=(AB-D)x where $A\in\mathbb{R}^{n\times r}$, $B\in\mathbb{R}^{r\times n}$, $D\in\mathbb{R}^{n\times n}$. Clearly $AB$ is a redundant parameterization of a rank $r$ matrix, but we use this notation to contrast the non-linear setting in Section 4.2. To ensure the drift is Hurwitz and the SDE has a stationary distribution, we assume $\|AB\|\leq\gamma<1$ and $D\succeq I$. Intuitively, $AB$ drive the dynamics while $D$ is a decay term to prevent unbounded dynamics. This parameterization is similar to the one used in Rohbeck et al. (2024), which also uses linear SDEs but considers hard interventions with having entirely new rows in the drift matrix, rather than shift interventions.
@@
Additionally, for theoretical tractability, we presume knowledge of the decay term $D$. This is a stronger assumption, but plausible in some applied contexts. For example, in single-cell genomics the decay rate of genes can be estimated with external experiments (e.g., from BRIC-seq (Imamachi et al., 2014)).
@@
If diagonal decay term $D$ is not inferred in advance, learning it simultaneously is comparable to the setting of robust PCA (Candès et al., 2011) or recovery of a diagonal plus a positive semidefinite low rank term (Saunderson et al., 2012). However, unlike the usual matrix completion setting where entries of the matrix are revealed uniformly at random, we have access to correlated low-rank measurements of the form $e_{i}c_{j}^{T}$ for $\{e_{i}\}_{i=1}^{n}$ the elementary basis. This setting is well studied with sub-Gaussian measurement vectors (Zhong et al., 2015), but the tools do not readily apply to our setting with non-random measurements and the constraints of the Lyapunov equation. Nevertheless, we observe in Section 5 that identifiability with an unknown diagonal decay matrix $D$ is empirically achieved, while still subject to the lower bound established in Theorem 4.4.
@@ 4.2 Nonlinear case
The proof is in Appendix A. Intuitively, the equilibrium points of the noiseless ODE approximate the mean of the SDE for small noise (note this is not true for larger noise, for example see Ma et al. (2015)). This suggests that in order to make use of the low-rank assumption on the dynamics, one must at least take advantage of second order moments of the SDE even in the small noise limit.
@@
The proof is given in Appendix A. This is one instance of perturbation theory for SDEs (Gardiner, 2021; Sanz-Alonso & Stuart, 2016). Equipped with this fact, one can consider access to the stationary distribution of an SDE, specifically the first and second-order moments, and inspect what happens now as the noise goes to zero. In this case, we obtain a non-trivial identifiability bound:
@@
We hypothesize the constraint that the SDE drift have a single absorbing point is unnecessary. If the ODE given by $\frac{dx}{dt}=v(x)+c_{i}$ had multiple stable equilibria, then in the limit of small noise, the SDE stationary distribution should approach a Gaussian mixture centered around each equilibrium point. The proof only requires certain linear independence conditions of vectors induced by each fixed point, so one intervention with multiple equilibria would offer essentially the same information as additional unimodal interventions. This is consistent with the observations in Lorch et al. (2024) that nonlinear SDE models generalize well with a limited number of interventions.
@@ 5 Experiments
To validate the theory given above, we consider experiments demonstrating the recovery of SDE parameters and generalization to unseen interventions. To show the broad applicability of the theory, we consider multiple loss functions: a loss acting directly on the parameters in the linear case (Rohbeck et al., 2024), a kernelized Stein discrepancy (Barp et al., 2019), and rollout loss for neural SDEs (Kidger et al., 2021).
@@ 5.1 Linear SDE Recovery
Because we are fitting linear SDEs where the stationary distribution is Gaussian, we can choose a very simple loss that matches the population mean and covariance of each intervented distribution, to verify the claim of identifiability (Theorem 4.4). This loss is akin to the one used in the Bicycle method introduced in Rohbeck et al. (2024) that fits interventional linear SDEs. Assuming the noise scale $\epsilon$ is known and $L$ denotes the true drift, we fit the parameters $\hat{A}$, $\hat{B}$, and $\hat{D}$. Letting our estimate for the drift be denoted by $\hat{L}:=\hat{A}\hat{B}-\hat{D}$, the loss function matches the mean and covariance of each intervention:
@@
We confirm our theory: when the decay is known, r2r-2 interventions fails to grant identifiability with poor performance and extremely high variance, even with many independent initializations, while $r$ interventions clearly succeed. When the decay isn’t known, we use a modest oversampling of $r\log(n)$ interventions. This amount is informed by the literature on matrix completion, where $\Omega(rn\log(n))$ revealed matrix entries were proven to be necessary for a rank $r$ matrix recovery problem (Candès & Tao, 2010). We scale down by a factor of $n$ because in our setting each intervention yields the mean vector of the perturbed linear SDE in $\mathbb{R}^{n}$, hence $n$ measurements.
@@ 5.3 Synthetic Nonlinear SDE Generalization
For an evaluation of Theorem 4.8, we apply the insight that identifiability doesn’t require knowing $\sigma$ and consider learnable MLPs. Learnable activations (Goyal et al., 2019) have been proposed before (Apicella et al., 2019) but to our knowledge not previously applied to SDE parameterizations.
To assess generalizability, we consider a loss function proposed by Lorch et al. (2024): the kernel deviation from stationarity (KDS), or equivalently the kernelized Stein discrepancy of the stationary distribution (Barp et al., 2019). When the parametric drift is defined as $v_{\theta}(x)=A\sigma(Bx)-Dx$ and the target stationary distribution is $\mu$, the KDS is,
@@
where $k$ is a given kernel and $\mathcal{A}_{x}$ is the SDE generator acting on the variable $x$ (for more details see Lorch et al. (2024)). We use samples from the true stationary distribution $\mu$ to approximate these expectations, further training details are given in the Appendix. This loss is nevertheless unstable when using noise small enough to reach the limit proven in Theorem 4.8, and therefore we assess generalizability according to performance on unseen interventions.
We observed numerical instability in the Sinkhorn divergence without large entropic regularization (Cuturi, 2013). Therefore, we follow Zhang et al. (2023); Lorch et al. (2024) and evaluate with the mean squared error (MSE) between the true and predicted distribution. Further details are given in Section B.3. The results are given in Table 2. We observe a clear benefit for low noise SDEs, which gets smaller as the noise gets bigger and hence further away from our proven settings. This is consistent with our theory, and also intuitive: as the noise gets larger, the stationary distribution gets smoother and has less dependence on the intervention vectors. Nevertheless, in the low noise regime we see that learnable activations improve prediction on test interventions without overfitting.
@@ 5.4 Simulated GRN SDE Generalization
We use the PerturbODE model from Lin et al. (2025), similar to our setup and modified to handle a loss for parameterized SDEs (Kidger et al., 2021). We consider the inferred GRN as the matrix multiplication of the weight matrices in the MLP that parameterizes the drift, corresponding to a first-order Taylor approximation of the dynamics. GRN recovery can be measured as a classification task of individual edges. As a baseline, we also consider the Bicycle interventional linear SDE method (Rohbeck et al., 2024). To simulate a known GRN, we use the boolean ODE based-simulator BEELINE (Pratapa et al., 2020). BEELINE uses the provided GRN to parametrize an SDE in the space of mRNA expression and protein expression, with physically plausible regulation dynamics. More precise details are given in the Appendix in Section B.4.

### v4

@@ 1 Introduction
However, in many natural settings, data is modeled as following an SDE even if one does not have access to explicit trajectories. Studies of ecological systems focus on the long-term survival of multiple species modeled by the quasi-stationary state of SDEs with environmental factors as perturbations (hening2021stationary). The application of flow cytometry to protein signaling networks under perturbation (sachs2005causal) is destructive and yields protein quantification at one time point, modeled using the stationary distributions of linear SDEs in varando2020graphical.
One highly motivating application is single-cell genomic sequencing with high-throughput CRISPR perturbations. Biologists are often interested in inferring the gene regulatory network (GRN) that characterizes the dynamics of gene expression, informing which genes should be targeted for treatment (dixit2016perturb). But the destructive nature of sequencing makes it impossible to observe the trajectory of a single cell at multiple time-points, and in general it is difficult to obtain any time-series genomics data due to the high expense. Therefore, practitioners often only collect data at the end of an experiment, i.e., from the stationary distribution of the system.
Understanding the dynamics is essential for extrapolating to unseen settings, but noise and latent confounding makes it non-trivial to determine the true dynamics. Causal disentanglement aims to learn causal factors in spite of these confounders, mainly focusing on directed acyclic graph (DAG) based methods. To demonstrate these methods are well-founded, there is considerable effort devoted to understanding which models have identifiability guarantees (lachapelle2022disentanglement). However, these models suffer from inherent weakness, in particular 1) being unable to represent cycles or 2) approximate continuous-time dynamical models.
There has been renewed interest in modeling with stochastic differential equations (SDE) directly (peters2022causal). In the genomic context, there is precedent for this type of modeling to represent the so- called “Waddington landscape” (waddington2014strategy), the hypothetical energy surface of cells. Furthermore, SDEs are commonly used for simulating transcriptomic datasets from a given gene regulatory network (pratapa2020benchmarking; dibaeinia2020sergio).
As demonstrated in the context of diffusion models, SDEs are fully expressive in practice and can accurately generate observational data (song2021maximum). But to identify the true underlying SDE requires assumptions on the model. Foundational theoretical works on identifying dynamical systems typically learn from many trajectories or even the infinitesimal generator of the dynamics (hansen2014causal), leaving open the harder setting of observing only the stationary distribution.
@@ 2.2 Stochastic Differential Equations
We will only consider autonomous systems, i.e. where the drift and noise terms have no dependence on time $t$. We will enforce the weak conditions on the drift and noise to guarantee a unique stationary distribution (berglund2021long), with a density $p$ that satisfies the Fokker-Planck equation,
@@ 2.4 Interventional SDEs
We denote the concatenated intervention column vectors by the matrix $C\in\mathbb{R}^{n\times k}$. In the causal disentanglement literature, shift interventions typically give fewer guarantees (buchholz2024learning; squires2023linear). Therefore, we assume knowledge of the interventions:
@@ 3.1 Causal Representation
In terms of modeling causality (pearl2009causality), the most popular underlying model is the structural causal model (SCM), which characterizes the conditional distribution of a random variable under arbitrary intervention. Learning an SCM typically requires very strong assumptions such as sparsity (scholkopf2021toward), interventional data (lachapelle2022disentanglement), parametric assumptions (peters2014identifiability), among many other results. Sparsity is a very common theme in these models, though it may also be expressed in an assumption that the number of latent variables is small, i.e. a low-rank constraint (fang2023low).
@@ 3.2 Causal Disentanglement with Interventional Data
The bulk of the literature on causal disentanglement focuses on SCMs with an underlying DAG. Relevant to our work are results that assume access to multiple interventional environments, either acting directly on the observed variables (brouillard2020differentiable) or identifying a latent model under some distributional assumptions (lachapelle2022disentanglement; squires2023linear; buchholz2024learning).
Some of these works have addressed the crucial limitation of acyclicity (zheng2018dags; lee2019scaling; atanackovic2023dyngfn), but without necessarily incorporating dynamics. Many works require hard interventions, where an intervened variable is a function of exogenous noise, although some can handle soft interventions (zhang2024identifiability).
@@ 3.3 Dynamical System Methods
Previous work has considered modeling perturbations with dynamical systems (peters2022causal). One can prove identifiability of SDE parameters from the generator (hansen2014causal) or trajectories (guan2024identifying). Other methods work in our harder setting where observed data is drawn from the stationary distribution under an SDE, and match a learned SDE under numerous interventions. These works focus on linear drift and intervention-dependent parameters (rohbeck2024bicycle) or non-linear drift with shift interventions (lorch2024causal). Notably, neither paper gives theory to confirm if these models are identifiable.
There are also methods specific to a particular scientific domain. In genomics, some methods act on pseudotime, an inferred notion of time from cell states when very few “real” timepoints are available (wang2024regvelo; hossain2024biologically). Although harder to obtain due the destructive nature of RNA sequencing, genuine temporal data (even with very few timepoints) can also be modeled with the intent of extracting a GRN (lin2025interpretable).
The closest work to ours proves an identifiability bound for linear SDEs under a strong sparsity assumption (dettling2023identifiability). However, their result doesn’t consider any interventional data, and the exact pattern of sparsity in the drift matrix is assumed to be known a priori, which is rarely the case in applied problems of interest. Another closely related work is guan2024identifying, which can simultaneously infer the drift and diffusion. However, this work applies only in an easier setting that assumes linearity, doesn’t study interventions, and assumes multiple temporal marginals rather than just the stationary distribution.
@@ 4 Main Results
We consider two main parameterizations of the drift as linear or a two-layer neural network (an MLP) to verify when the model may be uniquely identified up to appropriate invariances. Prior work has considered sparsity in the linear setting only (dettling2023identifiability), but we focus on parameterizations that project to a low-dimensional space. In other words, although the ambient dimension $n$ may be large, we assume the hidden dimension $r$ is much smaller, at least $n>2r$, and ideally the number of interventions to uniquely identify the parameters should scale with $r$.
@@ 4.1 Linear case
We start with the linear case and consider the parameterization v(x)=(ABD)xv(x)=(AB-D)x where $A\in\mathbb{R}^{n\times r}$, $B\in\mathbb{R}^{r\times n}$, $D\in\mathbb{R}^{n\times n}$. Clearly $AB$ is a redundant parameterization of a rank $r$ matrix, but we use this notation to contrast the non-linear setting in Section 4.2. To ensure the drift is Hurwitz and the SDE has a stationary distribution, we assume $\|AB\|\leq\gamma<1$ and $D\succeq I$. Intuitively, $AB$ drive the dynamics while $D$ is a decay term to prevent unbounded dynamics. This parameterization is similar to the one used in rohbeck2024bicycle, which also uses linear SDEs but considers hard interventions with having entirely new rows in the drift matrix, rather than shift interventions.
@@
Additionally, for theoretical tractability, we presume knowledge of the decay term $D$. This is a stronger assumption, but plausible in some applied contexts. For example, in single-cell genomics the decay rate of genes can be estimated with external experiments (e.g., from BRIC-seq (imamachi2014bric)).
@@
If diagonal decay term $D$ is not inferred in advance, learning it simultaneously is comparable to the setting of robust PCA (candes2011robust) or recovery of a diagonal plus a positive semidefinite low rank term (saunderson2012diagonal). However, unlike the usual matrix completion setting where entries of the matrix are revealed uniformly at random, we have access to correlated low-rank measurements of the form $e_{i}c_{j}^{T}$ for $\{e_{i}\}_{i=1}^{n}$ the elementary basis. This setting is well studied with sub-Gaussian measurement vectors (zhong2015efficient), but the tools do not readily apply to our setting with non-random measurements and the constraints of the Lyapunov equation. Nevertheless, we observe in Section 5 that identifiability with an unknown diagonal decay matrix $D$ is empirically achieved, while still subject to the lower bound established in Theorem 4.4.
@@ 4.2 Nonlinear case
The proof is in Appendix A. Intuitively, the equilibrium points of the noiseless ODE approximate the mean of the SDE for small noise (note this is not true for larger noise, for example see ma2015complete). This suggests that in order to make use of the low-rank assumption on the dynamics, one must at least take advantage of second order moments of the SDE even in the small noise limit.
@@
The proof is given in Appendix A. This is one instance of perturbation theory for SDEs (gardiner2021elements; sanz2016gaussian). Equipped with this fact, one can consider access to the stationary distribution of an SDE, specifically the first and second-order moments, and inspect what happens now as the noise goes to zero. In this case, we obtain a non-trivial identifiability bound:
@@
We hypothesize the constraint that the SDE drift have a single absorbing point is unnecessary. If the ODE given by $\frac{dx}{dt}=v(x)+c_{i}$ had multiple stable equilibria, then in the limit of small noise, the SDE stationary distribution should approach a Gaussian mixture centered around each equilibrium point. The proof only requires certain linear independence conditions of vectors induced by each fixed point, so one intervention with multiple equilibria would offer essentially the same information as additional unimodal interventions. This is consistent with the observations in lorch2024causal that nonlinear SDE models generalize well with a limited number of interventions.
@@ 5 Experiments
To validate the theory given above, we consider experiments demonstrating the recovery of SDE parameters and generalization to unseen interventions. To show the broad applicability of the theory, we consider multiple loss functions: a loss acting directly on the parameters in the linear case (rohbeck2024bicycle), a kernelized Stein discrepancy (barp2019minimum), and rollout loss for neural SDEs (kidger2021neural).
@@ 5.1 Linear SDE Recovery
Because we are fitting linear SDEs where the stationary distribution is Gaussian, we can choose a very simple loss that matches the population mean and covariance of each intervented distribution, to verify the claim of identifiability (Theorem 4.4). This loss is akin to the one used in the Bicycle method introduced in rohbeck2024bicycle that fits interventional linear SDEs. Assuming the noise scale $\epsilon$ is known and $L$ denotes the true drift, we fit the parameters $\hat{A}$, $\hat{B}$, and $\hat{D}$. Letting our estimate for the drift be denoted by $\hat{L}:=\hat{A}\hat{B}-\hat{D}$, the loss function matches the mean and covariance of each intervention:
@@
We confirm our theory: when the decay is known, r2r-2 interventions fails to grant identifiability with poor performance and extremely high variance, even with many independent initializations, while $r$ interventions clearly succeed. When the decay isn’t known, we use a modest oversampling of $r\log(n)$ interventions. This amount is informed by the literature on matrix completion, where $\Omega(rn\log(n))$ revealed matrix entries were proven to be necessary for a rank $r$ matrix recovery problem (candes2010power). We scale down by a factor of $n$ because in our setting each intervention yields the mean vector of the perturbed linear SDE in $\mathbb{R}^{n}$, hence $n$ measurements.
@@ 5.3 Synthetic Nonlinear SDE Generalization
For an evaluation of Theorem 4.8, we apply the insight that identifiability doesn’t require knowing $\sigma$ and consider learnable MLPs. Learnable activations (goyal2019learning) have been proposed before (apicella2019simple) but to our knowledge not previously applied to SDE parameterizations.
To assess generalizability, we consider a loss function proposed by lorch2024causal: the kernel deviation from stationarity (KDS), or equivalently the kernelized Stein discrepancy of the stationary distribution (barp2019minimum). When the parametric drift is defined as $v_{\theta}(x)=A\sigma(Bx)-Dx$ and the target stationary distribution is $\mu$, the KDS is,
@@
where $k$ is a given kernel and $\mathcal{A}_{x}$ is the SDE generator acting on the variable $x$ (for more details see lorch2024causal). We use samples from the true stationary distribution $\mu$ to approximate these expectations, further training details are given in the Appendix. This loss is nevertheless unstable when using noise small enough to reach the limit proven in Theorem 4.8, and therefore we assess generalizability according to performance on unseen interventions.
We observed numerical instability in the Sinkhorn divergence without large entropic regularization (cuturi2013sinkhorn). Therefore, we follow zhang2023active; lorch2024causal and evaluate with the mean squared error (MSE) between the true and predicted distribution. Further details are given in Section B.3. The results are given in Table 2. We observe a clear benefit for low noise SDEs, which gets smaller as the noise gets bigger and hence further away from our proven settings. This is consistent with our theory, and also intuitive: as the noise gets larger, the stationary distribution gets smoother and has less dependence on the intervention vectors. Nevertheless, in the low noise regime we see that learnable activations improve prediction on test interventions without overfitting.
@@ 5.4 Simulated GRN SDE Generalization
We use the PerturbODE model from lin2025interpretable, similar to our setup and modified to handle a loss for parameterized SDEs (kidger2021neural). We consider the inferred GRN as the matrix multiplication of the weight matrices in the MLP that parameterizes the drift, corresponding to a first-order Taylor approximation of the dynamics. GRN recovery can be measured as a classification task of individual edges. As a baseline, we also consider the Bicycle interventional linear SDE method (rohbeck2024bicycle). To simulate a known GRN, we use the boolean ODE based-simulator BEELINE (pratapa2020benchmarking). BEELINE uses the provided GRN to parametrize an SDE in the space of mRNA expression and protein expression, with physically plausible regulation dynamics. More precise details are given in the Appendix in Section B.4.